OpenStudy (anonymous):

Solve the equation algebraically:

OpenStudy (anonymous):

\[e ^{2x}-5e^x+6=0\]

OpenStudy (anonymous):

Firstly, assign \[e^x\] a variable -- let's just say x, and then plug in that variable. From there, factor the quadratic equation.

OpenStudy (anonymous):

\[(e^x-3)(x^x-2)=0\] \[e^x=3\[ \[x=ln(3)\] or \[e^x=2\] \[x=ln(2)\]

OpenStudy (anonymous):

I'm lost now.

OpenStudy (anonymous):

Wait... I got it. Thanks (: